OpenStudy (anonymous):

Find the length of the curve r=1+cos(theta)

OpenStudy (anonymous):

i got it to 2*int from 0 to pi sqrt(2+2cos(theta)) d(theta). I do not understand how to simplify it to 2*int from 0 to pi sqrt((4(1+cos(2theta))/2) d(theta) I can get the rest

OpenStudy (anonymous):

r^2=1+2costheta+cos^2theta and dr/dtheta = -sin(theta).. dr/dtheta^2 = sin^2(theta).. right?

OpenStudy (anonymous):

thus L = int from 0 to 2pi sqrt(2+2cos(theta) dtheta

OpenStudy (anonymous):

sorry, my equation and draw buttons do not work today...

OpenStudy (anonymous):

r² + (r')²= (1 + cosθ)² = 1 + 2cosθ + cos²θ + (-sinθ)² = 1 + 2cosθ

OpenStudy (anonymous):

well, the thing is √(1 + 2cosθ) can not be integrated

OpenStudy (anonymous):

that would = 2+2cos(theta) since cos^2theta +sin^2theta =1, right?

OpenStudy (anonymous):

oops yes. But still, it can not be integrated

OpenStudy (anonymous):

how do I get it from sqrt(2+2cos(theta) to (4(1+cos(2theta))/2 ?

OpenStudy (anonymous):

sqrt(4(1+cos(2theta))/2 i mean.

OpenStudy (anonymous):

then they pull the sqrt4 out front.. and are left with sqrt[(1+cos)theta)/(2)], which they change to cos (theta/2) and then integrate.

OpenStudy (anonymous):

unfortunately my trig identity skills are poor.

OpenStudy (anonymous):

are you sure the book did that? The integrand can not be integrated, so even if there were a trig identity, wolfram alpha would have done that. But it say that it can not be done

OpenStudy (anonymous):

hmm.. thats what my sol manual .. and also Chegg says. Chegg did the simplification slightly different, but they both show the same answers.

OpenStudy (anonymous):

ok, let me check for some trig identity

OpenStudy (anonymous):

what did they have as answer?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

what is the limit of integration??

OpenStudy (anonymous):

0 to 2pi for the cardioid.. i did 2* 0 to pi.. since its symmetric

OpenStudy (anonymous):

can you take a snapshot of their work? I don't have microsoft word to read the file

OpenStudy (anonymous):

hmm what do u want me to paste teh screenshot to?

OpenStudy (anonymous):

that works

OpenStudy (anonymous):

btw, answer is indeed 8. But it was found by wolfram alpha

OpenStudy (anonymous):

i don't have a way besides pasting the screenshot into a word document. how else can i get the screenshot to you?

OpenStudy (anonymous):

its okay, I will ask my professor I guess. I am not good enough apparently to figure this out lol. hopefully this will not be an exam question haha

OpenStudy (anonymous):

I think i might found the trig identity

OpenStudy (anonymous):

hahah FOUND IT!! XDDD

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

ok, so √(2 + 2cosθ) = √(2) √(1 + cosθ) yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cosθ = cos²(x/2) - sin²(x/2) <--- this bad retriceis the key XDD

OpenStudy (anonymous):

they changed a.s.s to retrice O.O

OpenStudy (anonymous):

heheh

OpenStudy (anonymous):

anyway, under the square root, we have 1 + cos²(x/2) - sin²(x/2)

OpenStudy (anonymous):

but, 1 - sin²(x/2) = cos²(x/2)

OpenStudy (anonymous):

so, cos²(x/2) + cos²(x/2) = 2cos²(x/2)

OpenStudy (anonymous):

oooooooh dang!!!!!!

OpenStudy (anonymous):

so you have √(2) √[ 2 cos²(x/2) ] = 2 | cos(x/2) |

OpenStudy (anonymous):

2 ∫ |cos(x/2)| dx, from 0 to 2pi = 8 XDDD

OpenStudy (anonymous):

that's so awesome.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you're welcome :DDD

OpenStudy (anonymous):

that identity is not on my identity sheet ... lol guess i will have to add it.

OpenStudy (anonymous):

good idea lol